OpenStudy (anonymous):

6=x+2/3 .. step by step ?

OpenStudy (thread252):

here is a step by step calculation to your question by symbolab

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

6 = x+(2/3) Lets multiple by three both sides 18 = x + 2 Now lets subtract two to both sides 16 = x LETS CHECK: 6 = 16+(2/3) 18 = 16+2 16 = 16